OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

What is the slope of \(y=-x+6\)?

OpenStudy (anonymous):

-x

OpenStudy (anonymous):

The slope is the coefficient of the \(x\) term, not the \(x\) term itself.

OpenStudy (anonymous):

Then -1, I think

OpenStudy (anonymous):

Right. What's the slope of a perpendicular line relative to a given line? If I have a line with a slope of \(\dfrac{a}{b}\), what is the slope of a line perpendicular to this one?

OpenStudy (anonymous):

b/a

OpenStudy (jdoe0001):

\(\bf \cfrac{a}{{\color{blue}{ b}}}\qquad negative\to -\cfrac{a}{{\color{blue}{ b}}}\qquad reciprocal\to -\cfrac{{\color{blue}{ b}}}{a}\)

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

Okay, so the slope would in fact be \(-\dfrac{b}{a}\), not \(\dfrac{b}{a}\). So now if the slope of \(y=-x+6\) is -1, what's the slope of a perpendicular line?

OpenStudy (anonymous):

It would be -B / A, correct?

OpenStudy (anonymous):

If \(\dfrac{a}{b}=-1\), what's \(-\dfrac{b}{a}\)?

OpenStudy (anonymous):

1?

OpenStudy (jdoe0001):

\(\bf -1\to -\cfrac{1}{{\color{blue}{ 1}}}\qquad negative\to +\cfrac{1}{{\color{blue}{ 1}}}\qquad reciprocal\to +\cfrac{{\color{blue}{ 1}}}{1}\to 1\)

OpenStudy (anonymous):

Yes. Now, using the point-slope formula for a line, you would plug in the given point (-4,-2) and the slope 1: \[y-y_0=m(x-x_0)\]

OpenStudy (anonymous):

Would the answer be D?

OpenStudy (anonymous):

?